### v1

# Draw with Thought: Unleashing Multimodal Reasoning for Scientific Diagram Generation

Abstract
Scientific diagrams are vital tools for communicating structured knowledge across disciplines. However, they are often published as static raster images, losing symbolic semantics and limiting reuse. While Multimodal Large Language Models (MLLMs) offer a pathway to bridging vision and structure, existing methods lack semantic control and structural interpretability, especially on complex diagrams. We propose Draw with Thought (DwT), a training-free framework that guides MLLMs to reconstruct diagrams into editable mxGraph XML code through cognitively-grounded Chain-of-Thought reasoning. DwT enables interpretable and controllable outputs without model fine-tuning by dividing the task into two stages: Coarse-to-Fine Planning, which handles perceptual structuring and semantic specification, and Structure-Aware Code Generation, enhanced by format-guided refinement. To support evaluation, we release Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations. Extensive experiments across eight MLLMs show that our approach yields high-fidelity, semantically aligned, and structurally valid reconstructions, with human evaluations confirming strong alignment in both accuracy and visual aesthetics, offering a scalable solution for converting static visuals into executable representations and advancing machine understanding of scientific graphics.
Figure 1: Overview of our Draw with Thought framework.
1 Introduction
Scientific diagrams, such as model architectures, system workflows, and algorithmic flowcharts, are foundational to scholarly communication, providing compact visual representations of structured ideas [25]. However, most diagrams are published as rasterized PDF or PNG images [61], discarding their underlying symbolic structure and making them non-editable, non-executable, and difficult to interpret or reuse programmatically [12, 17, 52].
Figure 2: Overview of our Draw with Thought. The framework for scientific diagram generation processes input diagrams through Coarse-to-Fine Planning (perceptual structuring and semantic layout planning) followed by Structure-Aware Code Generation to produce editable, high-fidelity diagram representations.
Recovering structured representations from such diagrams entails mapping raw visual inputs to symbolic layouts comprising nodes, edges, spatial relations, and logical groupings [9]—a task that lies at the intersection of visual perception and structured abstraction. While Vision-Language Models (VLMs) have shown strong capabilities in aligning images with structured outputs, as demonstrated in captioning [31, 14], visual QA [47], and diagram layout prediction [43], their application to scientific diagram parsing remains nascent and largely limited to visually regular or low-complexity inputs. Existing studies have adopted various target formats that differ in expressiveness and abstraction level, such as SVG for vector rendering [44, 61], TikZ for LaTeX-integrated graphics [8, 9], and Python-based code for chart generation [55]. While SVG enables resolution-independent rendering, its reliance on low-level geometric primitives limits semantic interpretability [9]; TikZ supports precise layout control but introduces parsing challenges due to procedural syntax and structural variability [8, 9]; and although Python formats are effective for domain-specific plots, they lack the modularity and compositional flexibility needed to capture the heterogeneity of general-purpose scientific diagrams [55]. In contrast, XML-based formats such as mxGraph offer an explicit and editable graph abstraction, encoding diagrams as structured layouts with semantically grounded nodes, edges, and geometric constraints. This makes them particularly well-suited as target representations for recovering the symbolic structure of scientific diagrams from raw visual inputs [43].
In summary, existing methods for scientific diagram parsing face three key limitations: (1) reliance on shallow or domain-specific formats that restrict structural expressiveness [61, 44]; (2) modeling paradigms that focus on layout detection or retrieval, lacking symbolic reasoning and structural supervision [61]; and (3) assumptions of clean vector inputs or templates [55], which hinder generalization to real-world, heterogeneous, rasterized diagrams.
Figure 3: Qualitative Comparison Between DwT and State-of-the-Art MLLMs. We visualize the output results. We observe that our model can understand the layout and the content of the image well, achieving better performance.
To address these limitations, we present Draw with Thought, a training-free framework that leverages Multimodal Large Language Models (MLLMs) for symbolic reconstruction of scientific diagrams. Our approach draws inspiration from cognitive load theory [41], which suggests that breaking down complex tasks into manageable steps facilitates reasoning, and from structure mapping theory [22], which models analogy-making as a process of aligning relational structures. Building on such insights, we design a cognitively grounded Chain-of-Thought (CoT) prompting strategy that guides MLLMs through two key stages: Coarse-to-Fine Planning (Section 4.1), which includes Perceptual Structuring and Semantic Specification, and Structure-Aware Code Generation (Section 4.2), which incorporates iterative Format-Guided Refinement. These stages reflect how humans process diagrams—by sequentially recognizing visual elements, inferring their relationships, and assembling them into structured representations. Unlike prior works that rely on vision-only perception [55] or heuristic-based templates [16], our framework introduces a unified, interpretable reasoning pipeline compatible with diverse diagram styles and semantic structures. It supports high-fidelity generation of editable mxGraph XML code without any model fine-tuning. Our key contributions include:
  • We introduce Draw with Thought, which guides MLLMs to generate executable mxGraph representations from rasterized scientific diagrams via cognitively grounded chain-of-thought prompting: covering visual decomposition, relational inference, and symbolic code synthesis.
  • We construct Plot2XML, a benchmark consisting of 247 complex diagrams. It spans key domains such as computer vision, multimodal model, and broader AI systems, covering most of the key use cases for scientific diagrams.
  • Comprehensive experiments across multiple MLLMs demonstrate that our approach achieves high fidelity, semantic alignment, and easy editability—all without the need for model fine-tuning.
Table 1: Comparison of Scientific Diagram Datasets for Structural Understanding and Representation. We evaluate datasets across multiple dimensions. Full support ($\checkmark$) indicates comprehensive capability, partial support ($\checkmark$) indicates limited capability, and no support ($\times$) indicates absence of capability. Multimodality notation: T: Text-only, I: Image-only, I+T: Multi-modal with both images and text. Data authenticity reflects whether datasets contain real-world scientific diagrams from research publications and the complexity metric is expressed as the difficulty of data generation and evaluation.
Dataset Domain Characteristics Multimodality Format Complexity
Figure Type Data Authenticity T I I+T
SciGraphQA[34] Scientific Charts $\times$ $\checkmark$ $\times$ $\times$ Text $\times$
Design2Code[46] Website UI $\checkmark$ $\times$ $\checkmark$ $\times$ HTML $\checkmark$
SketchFig[9] Hand-drawn Diagrams $\checkmark$ $\times$ $\checkmark$ $\times$ TikZ $\times$
MMCode[32] Algorithm Diagrams $\times$ $\checkmark$ $\checkmark$ $\checkmark$ Code $\checkmark$
Plot2XML Complex Research Diagrams $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ XML $\checkmark$
2 Related Work
2.1 Image-to-Markup Conversion
Image-to-markup conversion supports a wide range of diagram types. Existing Studies rely on low-level vector primitives like SVG [44, 61], which ensure resolution independence but offer limited semantic detail. Others use Python-based chart-generation scripts including Plot2Code [55], PandasPlotBench [20] and Chart2Code [64], which excel in producing domain-specific plots yet struggle with heterogeneous diagrams [9]. Math-centric methods translate geometric figures or handwritten equations into  code [10, 53] and address symbolic inference, but often remain confined to specialized content. Besides, existing datasets largely fall short in covering key dimensions required for structured scientific diagram understanding. SciGraphQA [34] lacks visual inputs and complex structure, Design2Code [46] focuses on UI layouts with HTML rendering but lacks diagram-level abstraction, SketchFig [9] includes hand-drawn diagrams but provides limited layout semantics and MMCode [32] supports multimodal prompts but lacks authentic scientific content and XML-level structure. In contrast, XML-based structured markup language [43] under mxGraph standards [29], captures nodes, edges, and layout constraints in a unified form—an especially relevant feature for complex scientific diagrams that involve diverse graphical elements and rich relational structures. Consequently, to handle the intricacies of scientific diagram generation [25], our work adopts mxGraph to define the scientific diagram parsing task as an image-to-XML pipeline, enabling more expressive and editable representations.
2.2 Multimodal Large Language Models in Code Generation
Recent advances in Multimodal Large Language Models (MLLMs), such as Qwen-vl [6] and GPT-4 [28] demonstrate strong abilities to extract structured meaning from complex visual inputs, achieving notable results in image captioning [14, 31], visual question answering [47], and scene-level reasoning [21]. Meanwhile, code-oriented LLMs, including InCoder [19], StarCoder [33], and Code Llama [24], excel in generating executable programs across multiple languages. Existing studies that integrate visual cues with textual embeddings [61, 43] often focus on simpler diagrams or charts [55], where step-by-step reasoning involving Chain-of-Thought [54], self-refine [37, 16], self-reflect [13] reasoning-and-acting [62, 56]and plan-and-solve [51, 42] can enhance interpretability. However, many approaches rely on standardized domains [61] and struggle to capture the rich semantics of real-world scientific diagrams [9]. Finetuning models to generate lengthy code like XML [60], SVG [44, 61] or LaTeX [9] introduces substantial computational overhead, especially for small-parameter LLMs that must handle highly varied and domain-specific content. To overcome these challenges and facilitate XML-based code generation from complex scientific images, our work adopts a training-free strategy that leverages MLLMs for robust spatial reasoning, domain-specific logic, and knowledge-intensive tasks.
Inspired by cognitive load theory [41], which advocates breaking intricate problems into manageable steps-and structure mapping theory [22]—which models human reasoning through the alignment of relational structures—we propose Draw with Thought, a framework that guides MLLMs to parse and synthesize diagrammatic information without the need for expensive model finetuning.
3 Plot2XML: A Benchmark for Scientific Diagram-to-Code Generation
As shown in Table 1, prior datasets fall short in authenticity or structure, while Plot2XML provides real scientific diagrams with verified XML annotations.
3.1 Data Collection and Curation
We introduce Plot2XML, a novel benchmark dataset consisting of 247 scientific diagrams carefully selected from influential conference papers across multiple domains. The collection spans model architecture diagrams, flowcharts, system pipelines, and conceptual frameworks essential for scientific communication. Unlike SVG representations[15, 48], our Plot2XML diagrams in mxGraph format exhibit a geometric increase in both element count and code complexity, making this the first large-scale collection of scientific diagrams with manually verified markup-based structural annotations. We sampled diagrams from top-tier conferences, ensuring temporal and venue diversity that reflects contemporary visualization practices in research. Further details about the dataset construction and annotation process are provided in Appendix A.
3.2 Annotation Process
Each diagram was meticulously annotated with its corresponding mxGraph XML representation through a rigorous crowdsourcing process [50]. A team of expert annotators manually recreated each diagram using Draw.io (diagrams.net). The process involved analyzing the original raster image, identifying all visual elements and connections, recreating the diagram with high fidelity, exporting the structured XML representation, and verifying quality through secondary review. This manual recreation approach ensures perfect alignment between visual diagrams and their XML code representations, creating reliable image-code pairs critical for training and evaluating diagram understanding models.
3.3 Complexity Analysis
Figure 4: Complexity analysis of scientific diagrams. Visualized as a radar chart over five dimensions: Connection Complexity, Graphical Complexity, Color Effects, Text Annotation, and Special Elements. The chart compares three difficulty levels—Easy (blue, 85 diagrams, 34.4%), Medium (coral, 98 diagrams, 39.7%), and Hard (green, 64 diagrams, 25.9%)—highlighting distinct structural patterns across complexity axes
To systematically categorize the diagrams, we developed a multi-dimensional complexity analysis framework evaluating each diagram across five key dimensions: Connection Complexity (relationships between elements), Graphical Complexity (visual sophistication), Color Effects (semantic encoding), Text Annotation (textual density), and Special Elements (non-standard visual components). Based on these dimensions, we classified diagrams into three difficulty levels. As shown in Figure 4, Hard diagrams exhibit consistently higher complexity scores across all dimensions, with particularly pronounced differences in Connection Complexity ($score=4.1$) and Graphical Complexity ($score=4.3$). Further details about the dataset analysis are provided in Appendix A.
Algorithm 1 Draw with Thought (DwT)
1:Rasterized scientific diagram $D_{scientific}$
2:Structured code $Y_{mxGraph}$ in mxGraph XML
3:Stage I: Coarse-to-Fine Planning for Structured Visual Understanding $\triangleright$ (Sec. 4.1)
4:Phase 1.1: Perceptual Structuring
5:$T_{percept}=\mathcal{M}_{MLLM}(D_{scientific},{Thought}_{percept})$
6:  $=(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector})$
7$\triangleright$ $T_{gestalt}$: perceptual grouping via Gestalt principles
8$\triangleright$ $T_{hierarchy}$: decomposition into visual primitives
9$\triangleright$ $T_{encoding}$: visual-to-semantic variable mapping
10$\triangleright$ $T_{connector}$: element linkage and routing topology
11:Phase 1.2: SemanticSpecification
12:$T_{hierarchy}=\mathcal{M}_{MLLM}(T_{percept},{Thought}_{hierarchy})$
13:  $=(\mathcal{R},\mathcal{E},\mathcal{L})$
$\triangleright \mathcal{R}$: semantic regions (e.g., Input, Output)
$\triangleright \mathcal{E}$: typed elements with class labels (e.g., Entity)
$\triangleright \mathcal{L}$: layout constraints (alignment, layering, connectivity)
17:Stage II: Structure-Aware Code Generation via Progressive Realization $\triangleright$ (Sec. 4.2)
18:Phase 2.1: Initial Structured Code Generation
19:$Y_{mxGraph}=\mathcal{M}_{MLLM}(T_{hierarchy},{Thought}_{\text{code}})$
20:  $=\bigcup_{i}Y_{i}$ =$(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{\text{edge}})$
21$\triangleright$ $Y_{\text{doc}}$: root declarations (mxfile, diagram)
22$\triangleright$ $Y_{\text{style}}$: visual style tokens (colors, borders, fonts)
23$\triangleright$ $Y_{\text{node}}$: nodes from $\mathcal{E}$ with geometry and IDs
24$\triangleright$ $Y_{\text{layout}}$: layout constraints from $\mathcal{L}$
25$\triangleright$ $Y_{\text{edge}}$: directed connectors with routing logic
26:Phase 2.2: Multi-Round Format-Guided XML Refinement
27:for $t=1$ to $T_{\text{refine}}$ do
28:     $Y_{mxGraph}^{(t)}=\mathcal{M}_{MLLM}(Y_{mxGraph}^{(t{-}1)},$
29:            ${RefineThought}^{(t)}_{format})$
30:     if Draw.ioVerifier($Y_{mxGraph}^{(t)}$) == valid then break
31:end for
32$\triangleright$ Verification ensures XML schema validity, rendering success, and structural consistency
33:return $Y_{mxGraph}^{(T^{*})}$
34$\triangleright$ $T^{*}\leq T_{\text{refine}}$, earliest valid XML accepted by Draw.io
4 Draw with Thought
Following [51], We propose Draw with Thought (DwT), a planning-driven framework that transforms a rasterized scientific diagram $D_{scientific}$ into structured, editable graph code $Y_{mxGraph}$ in mxGraph XML format by guiding a Multimodal Large Language Model $\mathcal{M}_{MLLM}$ through planning-and-generate Chain of Thought. Inspired by how humans decompose diagrams through sequential reasoning [41, 22]—from perceptual recognition to semantic understanding and symbolic execution—we explicitly separate the diagram-to-code task into two interpretable stages  [^1] Please see Appendix C.1 for prompt details.: (1) Coarse-to-Fine Planning involving ${Thought}_{percept}$ and ${Thought}_{hierarchy}$ (Section 4.1). (2) Structure-Aware Generation involving ${Thought}_{code}$ and ${RefineThought}_{format}$ (Section 4.2), as detailed in Algorithm 1 and Figure 2. This separation enables the model to reason over symbolic abstractions before emitting graphical code.
4.1 Coarse-to-Fine Planning for Structured Visual Understanding
Perceptual Structuring through Symbolic Abstraction.
To reduce perceptual complexity, we first induce a symbolic abstraction $T_{percept}$ that encodes diagram structure through layout grouping, hierarchy, visual encoding, and topology extraction. Inspired by GoT [16], we prompt Multimodal Large language Model $\mathcal{M}_{MLLM}$ to simulate perceptual reasoning via ${Thought}_{perceptual}$:
$T_{percept}\sim\mathcal{M}_{MLLM}(\cdot\mid D_{scientific},{Thought}_{percept})$ (1)
We decompose the output into four complementary components:
$T_{percept}=(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector})$
where $T_{gestalt}$ encodes perceptual grouping using Gestalt principles (proximity, similarity, continuity, closure), $T_{hierarchy}$ defines recursive decomposition of visual objects into primitives and their nesting, $T_{encoding}$ identifies mappings from visual variables to semantic roles (e.g., color $\mapsto$ category, size $\mapsto$ quantity), and $T_{connector}$ captures all connective topology between elements, including directionality and routing.
Semantic Specification through Layout-Aware Abstraction
Once perceptual reasoning is complete, the MLLM transitions to semantic abstraction, converting low-level visual cues into structured symbolic representations. Given $T_{percept}$, the model is guided by ${Thought}_{hierarchy}$ to define a semantically meaningful layout plan:
$T_{hierarchy}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{percept},{Thought}_{hierarchy})$ (2)
We represent the layout plan as:
$T_{hierarchy}=(\mathcal{R},\mathcal{E},\mathcal{L})$ (3)
where $\mathcal{R}$ denotes semantic regions (e.g., Input, Output), $\mathcal{E}$ is the catalog of typed visual elements with class labels and graphical roles (e.g., Process, Decision, Entity), and $\mathcal{L}$ formalizes spatial constraints among elements, defined as:
=$\mathcal{L}=\{\texttt{align}(e_{i},e_{j}),\ \texttt{connect}(e_{i}\rightarrow e_{j}),\ \texttt{layer}(e_{k},z)\}$ (4)
This abstraction step lifts diagram content from geometric perception to symbolic understanding. Importantly, the layout constraints in $\mathcal{L}$ preserve both the visual integrity and semantic logic of the diagram, serving as the backbone for code synthesis in the next stage.
4.2 Structure-Aware Code Generation via Progressive Realization
Initial Structured Code Generation.
Conditioned on $T_{hierarchy}$, we prompt $\mathcal{M}_{MLLM}$ to progressively synthesize the final XML code $Y_{mxGraph}$:
$Y_{mxGraph}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{hierarchy},{Thought}_{code})$ (5)
We decompose the XML structure into five submodules:
$Y^{(0)}_{mxGraph}=(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge})$ (6)
where $Y_{doc}$ defines root-level declarations and metadata containers (e.g., mxfile, diagram, mxGraphModel), $Y_{style}$ encodes style dictionaries with reusable keys, $Y_{node}$ instantiates shape primitives from $\mathcal{E}$ with geometry and identifiers, $Y_{layout}$ applies $\mathcal{L}$ constraints to compute coordinates and alignment, and $Y_{edge}$ constructs directional connectors with routing points, anchors, and labels.
Finally, we define the complete output as:
$Y^{(0)}_{mxGraph}$ $=\bigcup_{i}Y_{i},$ (7)
for $\quad\text{for }Y_{i}\in\{Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge}\}$
Each component serves a distinct role: $Y_{doc}$ initializes the document tree (e.g., mxfile, diagram); $Y_{style}$ defines reusable style tokens (e.g., color, font, stroke); $Y_{node}$ instantiates entities from $\mathcal{E}$ with geometry and ID; $Y_{layout}$ applies spatial constraints from $\mathcal{L}$; and $Y_{edge}$ encodes directed connectors with routing semantics.
Multi-Round Format-Guided XML Refinement.
While structurally grounded, initial outputs often contain XML formatting flaws—such as unclosed tags, incorrect nesting, or inconsistent styles for long-context for output during inference [7]. To ensure syntactic correctness and structural executability, we introduce a multi-round refinement module guided by format-level reasoning.
At each refinement step $t$, $\mathcal{M}_{MLLM}$ revisits the prior output $Y^{(t-1)}_{mxGraph}$ and produces an updated version using format-specific reasoning via ${RefineThought}^{(t)}_{format}$ prompting:
$Y^{(t)}_{mxGraph}\sim$ $\mathcal{M}_{MLLM}\big{(}Y^{(t-1)}_{mxGraph},\ {RefineThought}^{(t)}_{format}\big{)},$
$\quad t=1,\dots,T_{refine}$ (8)
Each refinement round injects XML schema constraints, tag correction heuristics, and layout logic. The refinement process is terminated early if the intermediate output passes verification in Draw.io:
if $\text{if }\texttt{Draw.ioVerifier}(Y^{(t)}_{\text{mxGraph}})=\texttt{valid}\text{ then break}$ (9)
This ensures that the final output $Y^{(T^{*})}_{\text{mxGraph}}$ is well-structured and executable. The resulting XML can be directly used in platforms such as draw.io, enabling downstream applications and interactive editing.
[…]
CLIP $\uparrow$ DINO $\uparrow$ FID $\downarrow$ Aesth. $\uparrow$ Tok. (K) CLIP $\uparrow$ DINO $\uparrow$ FID $\downarrow$ Aesth. $\uparrow$ Tok. (K) CLIP $\uparrow$ DINO $\uparrow$ FID $\downarrow$ Aesth. $\uparrow$ Tok. (K)
[…]
GPT-4o 0.72 0.87 172 3.95 1.5 0.51 0.69 208 3.81 2.0 0.38 0.58 243 3.66 2.9
Llama-3.2V-11B 0.73 0.88 161 4.17 2.2 0.54 0.75 199 4.02 3.1 0.44 0.64 213 3.71 3.6
Gemini 2.0 0.76 0.89 146 4.32 3.0 0.59 0.81 181 4.06 6.0 0.48 0.73 223 3.85 10.5
Grok 3 0.74 0.90 155 4.27 3.3 0.60 0.78 178 4.17 5.8 0.50 0.70 211 4.05 9.6
GPT-4.5 0.80 0.94 148 4.51 2.0 0.64 0.79 165 4.21 4.5 0.54 0.72 185 3.95 7.6
Qwen 2.5 VL-32B 0.75 0.88 158 4.49 1.9 0.62 0.81 168 4.34 3.6 0.52 0.73 195 4.17 6.6
Claude 3.5-sonnet 0.78 0.92 119 4.82 1.8 0.61 0.84 148 4.41 4.1 0.52 0.77 180 4.15 8.8
Claude 3.7-sonnet 0.82 0.93 73 4.85 2.0 0.68 0.86 112 4.50 8.9 0.60 0.78 150 4.25 14.0
[…]
Qwen 2.5 VL-32B+CoT 0.78 0.88 133 4.71 2.0 0.64 0.82 163 4.39 3.9 0.55 0.75 178 4.15 7.4
Claude 3.7-sonnet+CoT 0.83 0.94 55 4.88 3.3 0.71 0.88 98 4.52 9.9 0.63 0.81 139 4.28 16.9
[…]
Qwen 2.5 VL-32B+DwT 0.81 0.93 67 4.80 2.1 0.70 0.84 129 4.56 6.0 0.62 0.78 152 4.30 10.9
Claude 3.7-sonnet+DwT 0.87 0.95 39 5.12 3.6 0.75 0.91 57 4.93 11.2 0.70 0.86 85 4.72 24.9
5 Experiment
5.1 Experimental Setup
To provide a comprehensive evaluation, we compared Draw-with-Thought (DwT) against eight state-of-the-art MLLMs— GPT-4o  [^2] We use the ”gpt4o-2024-07-18” version of API for reproducibility. [39], Claude3.5-sonnet [4], Claude3.7-sonnet [5], Gemini-2.0 [23], Grok-3 [59], Llama-3.2V-11B[2] and Qwen2.5-VL-32B [6].
Following [44, 61], we evaluate performance across two key dimensions using four metrics: (1) Similarity, measured by CLIP Score [26, 47], DINO Score [11, 40], and FID [27], to assess semantic and visual alignment with ground truth diagrams; (2) Aesthetics, assessed using the Aesthetic Score [45], which reflects perceived visual quality. Representative results are shown in Figure 3, with full visualizations and quantitative comparisons provided in Appendix C.2.
Overview.
We conducted comprehensive experiments evaluating our Draw with Thought (DwT) method against state-of-the-art MLMs on our Plot2XML benchmark across three difficulty levels. Our experimental results in Table 2 reveal profound insights into the capabilities of MLLMs for scientific diagram reconstruction—a task that demands sophisticated visual-spatial reasoning, structural abstraction, and symbolic translation. Notably, our method consistently outperforms all baseline models across all metrics, with improvements of 10-20% on semantic alignment scores and up to 40% on visual quality metrics for the most challenging tasks. The performance patterns across models and difficulty levels illuminate the fundamental cognitive challenges in transforming rasterized diagrams into structured representations.
What drives our DwT superior performance?
Our method achieves remarkable improvements, particularly in Hard difficulty tasks, by fundamentally aligning with the cognitive processes underlying diagram comprehension[1, 3]. By guiding models through visual decomposition, structural inference, and symbolic code synthesis, our approach directly addresses the core challenges in scientific diagram parsing: maintaining coherent spatial-relational representations while translating between visual and symbolic modalities [58]. The exceptional improvement in Hard tasks demonstrates that our method effectively scaffolds the complex reasoning required to handle diagrams with intricate spatial arrangements, nested hierarchies, and complex logical relationships—precisely the elements that make scientific diagrams challenging to reconstruct.
How does complexity impact performance?
The consistent performance degradation observed across all models as task difficulty increases directly reflects the escalating cognitive demands of complex diagram parsing. This degradation is particularly pronounced in base models without reasoning scaffolds, aligning with investigations of MLLM cognitive and memory mechanisms [41], which indicate that complex visual-spatial tasks can overwhelm working memory when approached holistically [49, 36]. The disproportionate decline in FID scores compared to semantic alignment metrics suggests that models primarily struggle with maintaining spatial coherence and structural relationships rather than conceptual understanding.
What are current MLLMs’ fundamental limitations?
Current MLLMs face a fundamental bottleneck in scientific diagram reconstruction: capturing multi-level abstractions often requires generating outputs exceeding the model’s token limit, leading to truncation or content loss. While models like Claude 3.7-sonnet benefits from strong cross-modal consistency, its sharp performance drop on complex diagrams—reflected by a 26.8% decline in CLIP score and a 105.5% increase in FID from Easy to Hard tasks (Table 2)—reveals the inadequacy of implicit spatial reasoning [63]. Conversely, the 47.2% CLIP score drop observed in GPT-4o suggests that certain architectures fundamentally lack the capacity to preserve coherent spatial-relational representations across varying levels of diagrammatic complexity. Given that spatial arrangements encode underlying logical relationships, the inability to maintain such structure compromises semantic fidelity and highlights spatial-relational understanding as a critical yet underdeveloped competency in existing multimodal models.
Figure 5: Comparison of diagram reconstruction results across different approaches. (a) Raw image shows the original scientific diagram with a two-step process structure. (b) Claude-3.7-sonnet’s direct generation attempt shows reasonable structure but misses some key connections and node relationships. (c) Our Draw with Thought (DwT) approach produces a more accurate reconstruction that better preserves the hierarchical structure, node positions, and connection relationships from the original diagram.
5.2 Ablation Study
Our Draw with Thought consists of three key components:—Perceptual Structuring, Semantic Layout Planning, and Structured Code Generation—which collectively support the extraction, organization, and formalization of visual content from scientific diagrams. To further enhance structural reliability, our framework incorporates a hierarchical XML generation module coupled with a self-refinement mechanism that iteratively corrects syntactic inconsistencies.
As shown in Table 3, we conduct ablation studies on our Plot2XML using LLaMA 3.2-V-11B, evaluating CLIP and DINO for semantic alignment, aesthetic score for visual quality, and XML Validation for structural correctness (Table 3). The full model yields the best results (CLIP 0.657, DINO 0.815, Aesthetics 3.807, Validation 89%). Removing Perceptual Structuring reduces CLIP and DINO by 7.2% and 9.0%, respectively, indicating weaker semantic grounding, while excluding Layout Planning leads to the largest overall decline—CLIP drops 25.1%, aesthetics 7.0%, and validation falls to 73%—highlighting its role in spatial-structural coherence. Excluding Context Style Design causes moderate losses in CLIP (–11.4%) and aesthetics (–7.0%), suggesting its effect is primarily stylistic. Removing Hierarchical XML generation lowers validation to 66%, showing flat generation alone is insufficient, while refinement partially recovers it to 84%, underscoring the value of structural priors. These results demonstrate that each module contributes distinct functional value, and that high-fidelity, executable diagrams rely on the synergy of perceptual grounding, spatial reasoning, and structure-aware generation.
Table 3: Ablation experiment. Results for Draw with Thought and ablation with different components using LLaMA 3.2-V-11B on Plot2XML.
Framework Variant DINO CLIP Aesth. Valid.
LLama 3.2-V + DwT 0.815 0.657 3.807 89%
Component Ablations:
w/o Perceptual Structuring 0.742 0.610 3.726 85%
w/o Layout Planning 0.650 0.492 3.539 73%
w/o Context Style Design 0.795 0.582 3.541 87%
XML Generation Variants:
w/o Hierarchical XML 66%
Flat XML + Refine 84%
5.3 Human Evaluation
Following [30, 35, 18], we conducted a human evaluation to assess the alignment between automated metrics and human perception, combining absolute ratings and comparative judgments. For absolute ratings, domain experts used a 10-point scale [30] to evaluate each reconstructed diagram in terms of similarity to the original (1 = “completely different” to 10 = “identical”) and aesthetic quality (1 = “poor quality” to 10 = “professional quality”). For comparative evaluation, we employed Best-Worst Scaling (BWS) [35, 18], where annotators selected the best and worst outputs within a presented set. To manage annotation effort, comparisons were limited to the two strongest baseline models—Qwen2.5-VL and Claude 3.7-Sonnet—alongside our method (DWT), with human-annotated diagrams serving as reference.
As shown in Table 4, DWT achieves the highest scores in both similarity (7.3) and aesthetics (8.1), closely approaching the human reference (9.4 / 9.0), and outperforming Claude (6.1 / 7.0) and Qwen (5.6 / 6.3). In BWS, DWT receives the highest score (0.74), indicating consistent preference, while Qwen’s negative score (–0.87) reflects frequent disfavor. These results confirm DWT’s superior semantic fidelity, visual quality, and alignment with human preferences. Annotation protocol details for human evaluation are provided in Appendix B.
Table 4: Human Evaluation Results. DWT surpasses all automated methods and closely matches human-created diagrams across all dimensions.
Metric Qwen2.5V Claude-3.7 DWT Human
Similarity 5.6 6.1 7.3 9.4
Aesthetics 6.3 7.0 8.1 9.0
BWS Score -0.87 0.13 0.74 N/A
Correlation Analysis.
Spearman correlation between model metrics and human evaluations [38] shows that CLIP scores exhibit the strongest correlation with human judgments of similarity (Spearman’s $\rho=0.825$), followed by DINO ($\rho=0.763$) and FID ($\rho=-0.712$). This strong correlation with CLIP scores is particularly noteworthy, as it suggests that semantic alignment between the original and reconstructed diagrams—precisely what CLIP scores measures—is the most salient factor in human perception of reconstruction quality. The substantial negative correlation with FID scores indicates that while low-level visual fidelity matters, humans prioritize semantic and structural coherence over pixel-perfect reproduction when evaluating scientific diagrams.
6 Conclusion
In this paper, we present Draw with Thought, a training-free framework unleashing Multimodal Large Language Models (MLLMs) for reconstructing scientific diagrams from rasterized images into editable, executable mxGraph XML. Grounded in cognitive load [41] and structure mapping [22] theory, our Draw with Thought involving Coarse-to-Fine Planning for Structured Visual Understanding and Structure-Aware Code Generation via Progressive Realization—yields diagram representations that are both semantically faithful and structurally valid. To support systematic evaluation, we introduce Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations and multi-dimensional complexity analysis. We benchmarked our method across multiple state-of-the-art MLLMs, consistently demonstrating superior performance on high-complexity scientific diagrams through enhanced semantic alignment, layout fidelity, and XML validity.
In the future, we will further generalize our cognitively grounded pipeline to broader image-to-structure code generation tasks across scientific and industrial domains. Just as our training-free design avoids the prohibitive finetuning cost seen in [61], we also advocate for the development of efficient long-context adaptation strategies [57] for MLLMs—crucial for scaling symbolic reasoning under extended diagram code scenarios. Our Plot2XML benchmark, with its rich structural annotations and complexity taxonomy, offers a valuable testbed for evaluating such advances in context-efficient modeling and layout-aware fine-tuning.
Supplementary Material
Overview: In this supplementary material, we provide additional details and discussions corresponding to our proposed method Draw with Thought and benchmark Plot2XML. Specifically, it covers the following aspects:
  • In Appendix A, we describe in detail our multi-dimensional analysis framework used for constructing the Plot2XML dataset. It systematically categorizes scientific diagrams based on comprehensive complexity dimensions.
  • In Appendix B, we provide more details about the human evaluation protocol for assessing diagram reconstructions. It outlines a systematic process that utilizes expert evaluation to gauge the overall quality and visual fidelity of the generated diagrams.
  • In Appendix C, we provide full prompt details and additional case studies used in our Draw-with-Thought framework, detailing each stage of the pipeline to facilitate reproducibility and future research.
Table 5: Semantic Tokens Defined by Our Plot2XML.
Category Token Description
File Metadata [<|‘<mxfile>’>] Root element containing file metadata
[<|agent>] User agent string
[<|host|>] Host environment
[<|version|>] Version number used
[<|type|>] File type
Page Elements [<|‘<diagram>’|>] Defines a diagram page with a name and unique ID
[<|id|>] Unique identifier for the diagram/page
[<|name|>] Name of the diagram page
Graph Model [<|‘<mxGraphModel>’|>] Contains global graph settings and data
[<|dx|>] Horizontal offset
[<|dy|>] Vertical offset
[<|grid|>] Grid enable flag
[<|guides|>] Smart guides enabled
[<|connect|>] Allows cell connections
[<|arrows|>] Displays arrows on edges
[<|fold|>] Enables grouping folding
[<|pageWidth|>] Page width in pixels
[<|pageHeight|>] Page height in pixels
Document Structure [<|‘<root>’|>] Root container for all diagram cells
[<|background|>] Container for background layer elements
[<|components|>] Container for diagram components
[<|connectors|>] Container for connectors
[<|labels|>] Container for text labels and annotations
Graph Element [<|id|>] Unique identifier for the cell
[<|value|>] Display content (supports HTML formatting)
[<|style|>] Visual style string (colors, borders, rounding, etc.)
[<|parent|>] Identifier of the parent cell (defines hierarchy)
[<|edge|>] Flag indicating the cell is an edge
Geometry [<|‘<mxGeometry>’|>] Defines the position and size of a cell
[<|x|>] X coordinate
[<|y|>] Y coordinate
[<|width|>] Width in pixels
[<|height|>] Height in pixels
[<|relative|>] Relative positioning flag
Edge Control [<|‘<Array>’|>] Array container for edge control (bend) points
[<|‘<mxPoint>’|>] Defines a control point with x and y coordinates
Style Attributes [<|rounded|>] Enables rounded corners
[<|whiteSpace|>] Text wrapping setting
[<|html|>] Enables HTML rendering for text
[<|fillColor|>] Background fill color (hex value)
[<|strokeColor|>] Border/line color (hex value)
[<|strokeWidth|>] Border/line width (numeric)
[<|endArrow|>] Arrow style at the end of an edge
[<|curved|>] Edge curved flag
[<|edgeStyle|>] Edge drawing style
[<|entryX, entryY|>] Offset adjustments for target connection
[<|entryDx, entryDy|>] svg element attribute fy
[<|exitX, exitY|>] Relative source connection point ratios
[<|exitDx, exitDy|>] Offset adjustments for source connection
Text Styles [<|align|>] Horizontal text alignment (e.g., center)
[<|verticalAlign|>] Vertical text alignment (e.g., middle)
[<|fontStyle|>] Font style (numeric value for bold, etc.)
[<|fontSize |>] Font size (in points)
Appendix A Dataset Complexity Analysis Methodology
The annotation process for creating diagram-XML pairs followed a structured workflow where expert annotators examined each original diagram, reconstructed it using Draw.io (diagrams.net), exported the mxGraph XML representation, and verified quality through secondary review. This rigorous process ensured high-quality image-code pairs essential for training and evaluating diagram understanding models. Generated dataset employs a systematic approach to categorize scientific diagrams based on their visual and structural complexity. We developed a multi-dimensional analysis framework that evaluates each diagram across five key dimensions, each scored on a scale of 1-5, with higher scores indicating greater complexity.
A.1 Complexity Dimensions and Scoring Criteria
Graphical Complexity.
This dimension measures the visual sophistication of the diagram based on the number and variety of shapes and visual elements. We apply Canny edge detection to identify contours in each image and count significant contours (area > 10 pixels). The scoring is determined by the number of significant contours detected: images with fewer than 5 significant contours receive a score of 1; those with 5–14 receive a score of 2; 15–29 yield a score of 3; 30–49 yield a score of 4; and 50 or more receive the maximum score of 5. This metric effectively captures the visual density and structural complexity of the diagram.
Connection Complexity.
This dimension evaluates the intricacy of relationships between elements, including the number and arrangement of connecting lines and arrows. We employ Hough line transform to detect lines in the image after edge detection. The scoring follows a similar pattern to graphical complexity: images with fewer than 5 detected lines receive a score of 1; 5–14 lines receive a score of 2; 15–29 lines yield a score of 3; 30–49 lines yield a score of 4; and 50 or more lines receive a score of 5. This metric is particularly important for scientific diagrams where the relationships between components often convey critical information.
Special Elements.
[…]
Text Annotation.
This dimension assesses the density and complexity of textual elements, including labels and embedded explanations. We apply connected components analysis with Otsu thresholding to identify potential text regions, filtering components by size and aspect ratio. The number of identified text regions determines the score: images with fewer than 5 identified text regions receive a score of 1; 5–14 regions receive a score of 2; 15–29 regions receive a score of 3; 30–49 regions receive a score of 4; and 50 or more regions receive a score of 5. Text elements are crucial in scientific diagrams as they provide context and explanation, making this an important dimension for evaluation.
Color Effects.
This dimension evaluates the use of color for semantic encoding, highlighting, and visual organization. We analyze color diversity in HSV color space, calculating histograms for hue (18 bins) and saturation (10 bins), and counting significant color bins (containing > 1% of pixels). The combined count of significant hue and saturation bins determines the color diversity score: images with a a diversity value below 3 receive a score of 1 (grayscale images default to this score); 3–5 receive a score of 2; 6–9 receive a score of 3; 10–14 receive a score of 4; and 15 or more receive a score of 5. Color usage adds another layer of complexity to diagrams, often serving to group related elements or highlight important components.
A.2 Difficulty Level Classification
Based on the cumulative scores across all five dimensions, we classified the diagrams into three difficulty levels. The classification thresholds were determined using the 33.3% and 66.7% percentiles of the total complexity scores to ensure a balanced distribution across difficulty levels. This resulted in 85 diagrams (34.4%) classified as Easy with lower complexity scores, 98 diagrams (39.7%) as Medium with moderate complexity scores, and 64 diagrams (25.9%) as Hard with higher complexity scores.
This stratification enables more nuanced evaluation of diagram generation methods across varying levels of complexity. Easy diagrams typically feature simple structures with few elements and minimal color usage, making them suitable for baseline testing. Medium diagrams present moderate challenges with more complex structures and relationships. Hard diagrams represent the most challenging cases, often featuring intricate connections, specialized visual elements, and sophisticated color schemes that test the limits of current generation methods.
A.3 Implementation Details
The complexity analysis was implemented using OpenCV for image processing operations. For edge detection, we used Canny edge detection with thresholds of 100 and 200, which provided a good balance between detecting significant edges while filtering out noise. For line detection, we employed the probabilistic Hough transform with parameters: threshold=50, minLineLength=30, maxLineGap=10, which effectively identified meaningful line segments while ignoring minor discontinuities.
Connected components analysis used Otsu’s method for adaptive thresholding, which automatically determines optimal threshold values based on the image histogram. This approach proved robust across various image qualities and lighting conditions. To ensure consistent scoring across the dataset, all analysis was performed on images resized to a standard resolution. This standardization was necessary to prevent bias in the complexity metrics due to varying image sizes in the original diagrams.
The entire analysis pipeline was validated through manual inspection of a subset of diagrams across different complexity levels, confirming that the automated scores aligned with human perception of diagram complexity. This validation process helped refine the scoring thresholds and ensure the reliability of our complexity classification system.
Figure 6: We show the case of human drawings, including the scientific image and corresponding XML codes. Due to the limited space, we only demonstrate some parts of the XML code.
Appendix B Human Evaluation Protocol
To validate the alignment between automated metrics and human perception, we conducted a comprehensive human evaluation study assessing the quality of reconstructed diagrams. We developed a web-based evaluation interface (Figure 7) that presented evaluators with the original reference diagram alongside two generated versions for side-by-side comparison.
Figure 7: Web interface for human evaluation of diagram generation quality. The interface displays the original reference diagram (center) alongside two AI-generated versions (left and right). Evaluators assess each generated image on similarity to the original (how closely it matches the reference structure) and aesthetic quality (visual appeal).
B.1 Evaluation Design
We recruited 12 experts with experience in diagram design and evaluation, including researchers with advanced degrees in computer science and professionals with UI/UX design experience. From our test set, we randomly selected 50 diagram samples across various types.
For absolute ratings, participants used a 10-point scale to assess:
  • Similarity to original (1="completely different" to 10="identical"): evaluating node accuracy, connection relationships, structural preservation, and text content accuracy.
  • Aesthetic quality (1="poor quality" to 10="professional quality"): assessing layout balance, style consistency, element spacing, and visual appeal.
For comparative judgments, we employed Best-Worst Scaling (BWS), where participants simultaneously viewed the original diagram alongside three reconstructed versions (Qwen2.5VL, Claude3.7-sonnet, and human-created reference). They selected both the best and worst reconstructions, providing more sensitive relative comparisons and reducing rating biases.
B.2 Analysis Methods
We calculated mean scores and standard deviations for each model on similarity and aesthetic quality measures. For comparative judgments, we computed Best-Worst Scaling (BWS) scores using:
BWS score=times chosen besttimes chosen worsttotal $\text{BWS score}=\frac{\text{times chosen best}-\text{times chosen worst}}{\text{total evaluations}}$ (10)
To assess alignment between human perception and automated metrics, we calculated Spearman’s rank correlation coefficient ($\rho$) between human similarity ratings and various automatic evaluation metrics. This non-parametric measure quantifies the monotonic relationship between variables without assuming linearity:
$\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)}$ (11)
where $d_{i}$ represents the difference between ranks of human ratings and automated metric scores for each diagram, and $n$ is the number of evaluated samples. The resulting coefficients indicated moderate to strong correlations, suggesting our automated metrics reasonably approximate human judgment of diagram reconstruction quality.
XML Validity and Compilation.
A common challenge when generating mxGraph XML with LLMs is that the maximum token length may be insufficient to complete the XML code, resulting in compilation errors. We found that 87% of generated XML files were within context length and successfully compiled. For the remaining incomplete examples, we implemented a post-processing pipeline using the mxGraph JavaScript library to repair and complete the XML structure. This process involves parsing the incomplete XML, identifying missing closing tags, reconstructing the element hierarchy, and regenerating a valid XML document. In some cases, partial elements or attributes might be lost during this recovery process, particularly for complex diagrams with nested structures. However, this technical approach ensured that 100% of our generated XML files were valid and compilable, enabling comprehensive evaluation across our entire test set.
Appendix C Experiments Details for Draw-with-Thought
C.1 Prompt details for Draw-with-Thought
To support reproducibility, we release the prompts used in Draw-with-Thought, covering: Perceptual Structuring (Figure 8), Semantic Specification (Figure 9), Structure-Aware Code Generation (Figure 10), and Format-Guided XML Refinement (Figure 11).
Prompt ${Thought}_{perceptual}$ for Perceptual Structuring Instruction You are an expert in scientific visualization and draw.io (mxGraph XML) diagram design. Given a description or image of a scientific figure, your task is to deconstruct and reconstruct the visual structure using mxGraph XML format through a multi-stage analysis pipeline. 1. Gestalt Perception Analysis: In the realm of Gestalt perception, analyze the graph by identifying the relationship between graphical elements and the background, recording proximity-based groupings, examining similarity patterns among visual components, and noting the continuity in connector elements as well as any applicable closure patterns. In addition, record the overall symmetry and balance of the composition. 2. Hierarchical Decomposition:
Decompose the graph hierarchically by identifying its primary visual objects and their boundaries, breaking complex objects into basic mxGraph shapes, and establishing parent-child relationships among the elements. This involves recognizing containment and spatial nesting relationships as well as recording the visual layering (z-order) of the components.
3. Visual Encoding Analysis:
Conduct a visual encoding analysis to discern how data variables are mapped to visual attributes; document the color encoding scheme and its corresponding meaning; analyze the encoding of size, position, and orientation; and recognize the significance of different line styles, thicknesses, and text attributes.
4. Connector Relationship Mapping:
Analyze connector relationships by identifying every connection element (such as arrows and lines), recording the source and target for each connection, examining routing patterns and bend points, noting the specific styles and arrowhead designs used (including their meanings), and recognizing any labeled connections along with the positioning of their associated text.
Examples: Input: {Diagram $D_{Scientific}$} Answer: {$T_{\text{percept}}$ = ($T_{\text{gestalt}}$, $T_{\text{hierarchy}}$, $T_{\text{encoding}}$, $T_{\text{connector}}$)} Output Format: A structured representation of the graph reconstruction task as a four-tuple of analysis components: ($T_{\text{gestalt}}$, $T_{\text{hierarchy}}$, $T_{\text{encoding}}$, $T_{\text{connector}}$) $\triangleright$ $T_{\text{gestalt}}$: perceptual grouping via Gestalt principles $\triangleright$ $T_{\text{hierarchy}}$: decomposition into visual primitives and containment structure $\triangleright$ $T_{\text{encoding}}$: mapping of data variables to visual attributes $\triangleright$ $T_{\text{connector}}$: analysis of element linkage and connector routing topology
Figure 8: Prompt ${Thought}_{perceptual}$ for Perceptual Structuring
Prompt ${Thought}_{hierarchy}$ for Semantic Specification Instruction Based on the prior perception analysis, you are now tasked with constructing a complete semantic specification of all mxGraph elements and their associated styles required to reconstruct the scientific figure in exact XML form. This specification must eliminate all ambiguity in graphical representation, ensuring faithful one-to-one correspondence between perception and rendering. Step 1. Element and Shape Mapping: Identify the required shape libraries from draw.io (e.g., basic, UML, BPMN), and determine whether any custom shapes must be defined. Each identified visual component should be mapped to its corresponding mxGraph primitive, and any shape modifications or overrides must be recorded in full. Step 2. Style Definition and Custom Attributes: For each element type, define its visual style attributes with precision. Assign exact color values in HEX or RGB; specify fill types such as solid, gradient, or pattern; and configure border and connector line styles including width, dash pattern, and color. Text attributes should be detailed, including font family, size, color, alignment, and shadow effects where relevant. List any additional element-level attributes that may influence rendering or interaction, such as tooltips, conditional visibility, dynamic formatting, or metadata bindings. Include layer assignments for managing visual stacking or grouping semantics. Step 3. Connector Specification: For all connectors, specify routing type (e.g., straight, orthogonal, curved), arrowhead style and size, endpoint locations (including custom shape ports), and any routing constraints (such as fixed waypoints, edge snapping, or anchor preferences). Examples: Input: {$Diagram\ D_{Scientific}$ + $T_{percept}$}  Answer: {$T_{hierarchy}$ = ($S_{element}$, $S_{style}$, $S_{connector}$)} Output Format: A structured three-tuple describing the full reconstruction plan for mxGraph rendering: ($S_{\text{element}}$, $S_{style}$, $S_{connector}$) $\triangleright$ $S_{element}$: mapped shapes and required libraries for all components $\triangleright$ $S_{style}$: detailed visual styles for each element class $\triangleright$ $S_{connector}$: connector styles, endpoints, arrowheads, and routing logic
Figure 9: Prompt ${Thought}_{hierarchy}$ for Semantic Specification
Prompt ${Thought}_{code}$ for Structure-Aware Code Generation Instruction Based on the element specifications defined earlier, construct a complete base mxGraph XML document to replicate the scientific figure in draw.io. This implementation must conform precisely to the official XML schema used by draw.io, ensuring valid structure, correct style reference, and full compatibility with the latest version of the editor. Step 1. Document Root and Structural Initialization: Begin with the root <mxfile> element including appropriate namespace declarations and metadata. Create a <diagram> element with a unique id and human-readable name, followed by a well-formed <mxGraphModel> with attributes like dx, dy, grid, and gridSize. Within the model, ensure that the <root> contains an initial cell structure, including document-level configuration such as canvas bounds and page format settings. Step 2. Style Definitions: Define reusable visual styles via mxStylesheet. Implement all gradient fill settings, stroke colors, and border patterns for shape types. Configure connector styles including arrowheads, edge routing preferences (e.g., orthogonal, entity-relationship, or curved), and line styles. Specify text styles with font name, size, color, alignment, and other attributes like bold, italic, or shadow effects. Step 3. Layering and Grouping Logic: Define logical layers such as background, content, and annotation, setting their visibility, lock status, and z-index ordering. Within each layer, instantiate parent cells that represent structural groups, containers, or component modules as specified earlier. All shapes and connectors should be placed under the appropriate parent or layer node to preserve semantic organization. Step 4. Coordinate System and Layout Configuration: Set diagram origin, scale, and canvas boundaries. Define grid spacing, snapping alignment, and page layout. If applicable, configure ruler or guideline settings. Ensure that each graphical element is positioned with precise coordinate values relative to the canvas and scaling system. Examples: Input: {$Diagram\ D_{Scientific}$ + $T_{hierarchy}$}  Answer: {($Y_{\text{doc}}$, $Y_{\text{style}}$, $Y_{\text{node}}$, $Y_{\text{layout}}$, $Y_{\text{edge}}$)} Output Format: A complete mxGraph XML structure expressed as a five-tuple: ($Y_{doc}$, $Y_{style}$, $Y_{node}$, $Y_{layout}$, $Y_{edge}$) $\triangleright$ $Y_{doc}$: root declarations (<mxfile>, <diagram>) with namespace and metadata $\triangleright$ $Y_{\text{style}}$: visual style tokens for all elements (colors, borders, fonts, fill patterns) $\triangleright$ $Y_{node}$: diagram elements from $\mathcal{E}$ with unique IDs and geometry attributes $\triangleright$ $Y_{layout}$: layout constraints from $\mathcal{L}$ including alignment, nesting, layering $\triangleright$ $Y_{edge}$: directed connectors with anchor points, routing strategies, and arrowhead metadata
Figure 10: Prompt ${Thought}_{code}$ for Structure-Aware Code Generation
Prompt ${RefineThought}^{(t)}_{format}$ for Format-Guided XML Refinement. Instruction You are an expert in scientific diagram synthesis and mxGraph XML generation for draw.io diagrams. Your role is to process an existing mxGraph XML snippet (which may contain formatting or semantic issues) and return a corrected and fully valid XML output that adheres to the draw.io schema. [INPUT] XML Code: {PRESIGNED_XML_CODE} Error Message: ’Error loading file (non-drawing file (error on line 552 at column 10: StartTag: invalid element name))’ Step 1: SYNTAX_CHECK - Examine the XML code for syntax errors including missing/unclosed tags, improper nesting, or malformed attributes. - Focus on the reported error location for malformed tags or invalid element names. Step 2: SCHEMA_VALIDATION - Confirm that the structural elements <mxfile>, <diagram>, <mxGraphModel>, and <root> exist and are properly nested. - Verify compliance with the official draw.io schema. Step 3: HIERARCHY_INTEGRITY - Ensure nodes, edges, groups, and styles have correct parent-child relationships. - Validate containment structure and identifier linkage integrity. Step 4: CONNECTOR_AND_STYLE_CHECK - Review all edges for valid source/target references and appropriate connector styles. - Confirm that all style attributes (e.g., fillColor, strokeColor, fontSize) are well-formed and applied consistently. Step 5: OUTPUT_QUALITY - Output a corrected XML that is well-formed, semantically valid, and fully compatible with the draw.io editor. - Ensure the final layout and style faithfully reproduce the intended diagram structure. Do not return any comments, explanations, or intermediate results. Only output the final corrected mxGraph XML code.
Figure 11: Prompt ${RefineThought}^{(t)}_{format}$ for Format-Guided XML Refinement
C.2 Additional Case Studies for Draw-with-Thought
As depicted in Figure 12, Figure 13 & Figure 14, we present representative case studies that qualitatively illustrate the capabilities and limitations of our framework. These examples cover output visualization, human-level comparison, and XML code generation, offering deeper insights into how the model performs across varying levels of diagram complexity and structural demands.
Figure 12: More Qualitative Visualization Results of DwT and Plot2XML
Figure 13: Here we compare the Plot2XML output results with human drawings to explore the boundary of the model. When the images are complex, Plot2XML has an overall understanding of the content and layout of the image and can replicate the image to a certain extent, but there are still issues with some details.
Figure 14: Example XML Code Generation. Here We present an example of our model output, omitting some of the XML code due to limited space.